Solve for x
6x^2-2x-3=0

Answer by Ave(106)

use "Reverse FOIL" to factor
(2x+1)(3x-3) set each of those = 0
2x+1=0
2x =-1
x=-1/2
3x-3=0
3x=3
x=1
{-1/2, 1}